# itemsep=10pt,topsep=0pt,parsep=0pt,partopsep=0pt

### 03 Jan 2024 v2

@@ 3.2 Instance-level Evaluation for ReasonSeg
The ReasonSeg-Inst dataset is curated from samples of COCO [3] and ADE20K [39] databases. Based on the aforementioned process for prompting GPT-4V, finally, we have collected 62K image-query pairs for the training phase, and 1.3K for evaluation. It has come to our attention that the low-resolution images or tiny object instances within COCO and ADE20K can impede the GPT-4V’s ability to generate satisfactory instruction-tuning data. In light of this, we have adopted a filtration criterion, discarding the images whose size is below 512x512 and objects whose area is under 400 square pixels, to ensure the integrity of the dataset for effective instruction tuning data generation. More examples are shown in Figure 4 at the end of this manuscript.
@@ 4.1 Implementation Details
Training.   The training configuration for LISA++, such as training epochs, learning rate, losses and loss weights, etc., is predominantly in line with the framework of the original LISA model [9]. The notable divergence in LISA++ is the incorporation of new instruction tuning sets, specifically designed for refining reasoning in instance segmentation and SiD. To accomplish this, we utilize the COCO [3] and ADE20K [39] datasets. It is important to note that the images from COCO and ADE20K were already employed in the training of the original LISA model, hence no additional visual data is introduced for training LISA++.
Evaluation.   For evaluation on ReasonSeg-Sem, we follow the setting of the original LISA [9] where cIoU and gIoU are reported. As for ReasonSeg-Inst which requires instance segmentation, we report the standard COCO metrics including AP (averaged over IoU thresholds), AP50, and AP75 for the evaluation of ReasonSeg-Inst. AP is evaluating using mask IoU.
Metric Method
LISA-7B LISA-13B LISA++-7B
AP50 46.2 46.2 58.7
AP75 46.2 45.0 55.5
mAP 40.4 44.0 54.0
@@ 4.2 Results
In comparison to LISA-7b, our LISA++-7b achieves significant improvement in ReasonSeg-Inst scenario. Specifically, LISA++ achieves 58.7%, 55.5%, and 54.0% in AP50, AP75, and mAP, respectively. Even when compared to the more powerful LISA-13b, utilizing the 7b LISA++ still resulted in improvements of 12.5%, 10.5%, and 10.0% in AP50, AP75 and mAP, respectively. These results further demonstrate the effectiveness of our model in instance segmentation. We hope LISA++ to serve as the new baseline for the ReasonSeg-Inst task.

### 22 Jan 2024 v3

@@ 3.2 Instance-level Evaluation for ReasonSeg
The ReasonSeg-Inst dataset is curated from samples of COCO [3] and ADE20K [39] databases. Based on the aforementioned process for prompting GPT-4V, finally, we have collected 62K image-query pairs for the training phase, and 1.8K for evaluation. It has come to our attention that the low-resolution images or tiny object instances within COCO and ADE20K can impede the GPT-4V’s ability to generate satisfactory instruction-tuning data. In light of this, we have adopted a filtration criterion, discarding the images whose size is below 512x512 and objects whose area is under 400 square pixels, to ensure the integrity of the dataset for effective instruction tuning data generation. More examples are shown in Figure 7 at the end of this manuscript.
@@ 4.1 Implementation Details
Training.   The training configuration for LISA++, such as training epochs, learning rate, losses and loss weights, etc., is predominantly in line with the framework of the original LISA model [9]. The notable divergence in LISA++ is the incorporation of new instruction tuning sets, specifically designed for refining reasoning in instance segmentation and SiD. To accomplish this, we utilize the COCO [3] datasets. It is important to note that the images from COCO were already employed in the training of the original LISA model, hence no additional visual data is introduced for training LISA++.
Evaluation.   For evaluation on ReasonSeg-Sem, we follow the setting of the original LISA [9] where cIoU and gIoU are reported. As for ReasonSeg-Inst which requires instance segmentation, we report the standard COCO metrics including AP, AP50, AP75, AP-small, AP-medium and AP-large for the evaluation of ReasonSeg-Inst. AP is evaluating using mask IoU.
Metric Method LISA-7B LISA-13B LISA++-7B AP50 13.7 15.7 34.1 AP75 6.6 8.1 22.1 mAP 7.2 8.6 21.5 AP-small 0.9 1.0 2.5 AP-medium 3.3 5.0 16.7 AP-large 19.4 21.3 47.1
@@ 4.2 Results
In comparison to LISA-7b, our LISA++-7b achieves significant improvement in ReasonSeg-Inst scenario. Specifically, LISA++ achieves 34.1%, 22.1%, and 21.5% in AP50, AP75, and mAP, respectively. Even when compared to the more powerful LISA-13b, utilizing the 7b LISA++ still resulted in improvements of 18.4%, 14.3%, and 12.9% in AP50, AP75 and mAP, respectively. These results further demonstrate the effectiveness of our model in instance segmentation. We hope LISA++ to serve as the new baseline for the ReasonSeg-Inst task.
@@
Figure 4: Qualitative Results for Image Captioning with the Instance Segmentation.
Figure 5: Qualitative Results for Instance Segmentation.
4.3 Qualitative Results
To further demonstrate the effectiveness of our LISA++, we show the qualitative results in this section.
Image Captioning with Instance Segmentation  As shown in Figure 4, we show LISA++’s qualitative results for Image Captioning with the Instance Segmentation. LISA++ can yield more natural and flexible responses, by integrating segmentation results into textual responses, termed the ability of Segmentation during the natural captioning.
Instance Segmentation  Additionally, we assess the instance segmentation performance in Figure 5. These questions require advanced reasoning capabilities and pose challenges due to potential interference from other objects. Notably, our proposed LISA++ demonstrates exceptional accuracy in recognizing the queried objects and conducting precise segmentation. For instance, it accurately segments objects such as girls, white cars, and whole oranges.
Multi-turn Dialogue with Instance Segmentation  Furthermore, to conduct a more comprehensive evaluation of LISA++’s performance in the context of multi-turn dialogue with instance segmentation, we employ the ”SiD with InstSeg” instructions template and present the qualitative results in Figure 6. Our proposed LISA++ has the capability to gain a comprehensive understanding of the scene and carry out instance segmentation on the two individuals depicted in the image. Furthermore, LISA++ exhibits well perception of color attributes, such as ”white,” and spatial relationships, such as ”holding the bat.”
@@ 5 Conclusion
Figure 6: Qualitative Results for Multi-turn Dialogue with the Instance Segmentation.